OpenStudy (anonymous):

evaluate the expression for a=5 and b=2 -a+6b^2-9

OpenStudy (anonymous):

-5 + 6*(2^2) - 9 = 24 -5 - 9 =10

OpenStudy (aric200):

?

OpenStudy (aric200):

Did you figure it out or do you need help

OpenStudy (aric200):

He broke it down about as simple as it gets

OpenStudy (anonymous):

I believe he needed help, and I gave the solution. Correct me if I'm wrong @keenonhenderson

OpenStudy (anonymous):

\[a ^{3}-2b ^{2}-1 your \right\]